OpenStudy (anonymous):

I GIVE MEDALS! Circle P has a radius of 7 and Circle R has a diameter of 8. The circumference of Circle P is how much greater than the circumference of Circle R.

OpenStudy (aggie12341):

103.62

OpenStudy (anonymous):

Thanks! But what would that be in Pie? A. pie B. 6(pie) C. 8(pie) D. 16(pie) E.. 33(pie)

OpenStudy (campbell_st):

For circle P the circumference can be found using \[C = 2 \times \pi \times r\] just substitute your radius (r) and calculate for circle R you need \[C = \pi \times d\] substitute your diameter (d) and calculate then find the difference in the 2 measurements

OpenStudy (aggie12341):

d

OpenStudy (anonymous):

Thanks!

OpenStudy (aggie12341):

medal

OpenStudy (anonymous):

i gave you it